OpenStudy (anonymous):

The function H(t) = -16t2 + vt + s shows the height H(t), in feet, of a projectile launched vertically from s feet above the ground after t seconds. The initial speed of the projectile is v feet per second. Part C: Another object moves in the air along the path of g(t) = 31 + 32.2t where g(t) is the height, in feet of the object from the ground at time t seconds. Use a table to find the approximate solution to the equation H(t) = g(t), and explain what the solution represents in the context of the problem? [Use the function H(t) obtained in Part A, and estimate using integer values]

OpenStudy (anonymous):

1 only need help with parts C and D but here is the rest of it. Part A: The projectile was launched from a height of 96 feet with an initial velocity of 80 feet per second. Create an equation to find the time taken by the projectile to fall on the ground. Part B: What is the maximum height that the projectile will reach? Show your work. Part D: Do H(t) and g(t) intersect when the projectile is going up or down, and how do you know?

OpenStudy (anonymous):

@amistre64 @jagr2713 @shifuyanli

OpenStudy (anonymous):

Will give a medal.

OpenStudy (anonymous):

1t doesn't explain the question,

OpenStudy (anonymous):

hold on please

OpenStudy (anonymous):

can you please just explain how to do this.

OpenStudy (anonymous):

okay iwill try

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

mhm but if you want, you can solve it algebraically to find the solutions H(t) = g(t) −16t2+80t+96=31+32.2t −16t2+80t−32.2t+96−31=0 −16t2+47.8t+65=0 then use the quadratic formula to solve for the values of t t=−b±b2−4ac−−−−−−√2a where a=-16, b=47.8, c=65

OpenStudy (anonymous):

are the two interception points (-1,-1) and (4,160)?

OpenStudy (anonymous):

lets intersect at two points, since there's a solution when you set H(t)=g(t) and for the the value of "t" i got a different answer t is approximately 4.002 and -1.015

OpenStudy (anonymous):

1t said approximate so 1 rounded.

OpenStudy (anonymous):

mhm try looking at this link i found

OpenStudy (anonymous):

okay so the points are correct. How would 1 get part D?

OpenStudy (anonymous):

http://openstudy.com/study#/updates/53475411e4b01730eeaf944c i think

OpenStudy (anonymous):

okay that makes sense but these points would intersect when it was going up and down?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay thank you so much!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

1 may need more help later.

OpenStudy (anonymous):

okya tag me when you do